OpenStudy (anonymous):

JUST NEED CHECKED!! Suppose that x* means 5x^2 - x then what does (y+3)* mean? Do you plug in (y+3) for x, expand the squared, and the (y+3)'s cross out to 5(y+3)? (y+3)* = 5y+15

hero (hero):

Interesting...

hero (hero):

x* = 5x^2 - x y* = y+3 just a guess

OpenStudy (mertsj):

My thought would be that x* is like a function. f(x) = 5x^2-x x* = 5x^2-x (y+3)* = 5(y+3)^2-(y+3)

OpenStudy (anonymous):

I think so too. But wouldn't you expand it after that? 5(y+3)(y+3)-(y+3) Then the (y+3)s cross out leaving just 5(y+3)?

hero (hero):

Yeah, that probably is right. the y variable threw me off

OpenStudy (mertsj):

Yes. I think you should go ahead and square and then collect like terms.

OpenStudy (anonymous):

Okay, I think my final answer is 5y+15, since (y+3) are like terms.

OpenStudy (mertsj):

\[5(y+3)^{2}-(y+3)=5(y^2+6y+9)-y-3=5y^2+30y+45-y-3\]

OpenStudy (mertsj):

\[5y^2+29y+42\]

OpenStudy (anonymous):

So the last (y+3) doesn't cross out anything?

OpenStudy (anonymous):

Doesn't take a y out of the squared term?

OpenStudy (mertsj):

Remember the "powers first" rule in order of operations.

OpenStudy (mertsj):

This could be the approach you favor:

OpenStudy (mertsj):

\[5(y+3)^2-(y+3)=(y+3)[5(y+3)-1]=(y+3)[5y+15-1]=\]

OpenStudy (mertsj):

\[(y+3)(5y+14)=5y^2+14y+15y+42=5y^2+29y+42\]

OpenStudy (anonymous):

Thank you so much! You really helped. :)

OpenStudy (mertsj):

yw